# FinEval-KR: A Financial Domain Evaluation Framework for Large Language Models’ Knowledge and Reasoning

### v1

@@ A.2 Results Analysis
A.3 The Dataset and Prompts Used in Preliminary Experimemts
@@ Appendix C Data Sources for the FinEval-KR Dataset
When constructing a benchmark dataset to evaluate knowledge and reasoning abilities in finance, ensuring the authoritativeness and comprehensiveness of the data sources is paramount. To achieve this, we selected nine widely recognized and classic textbooks from various branches of modern finance. These cover subfield including corporate finance, investments, financial institutions and markets, money and banking, international finance, financial risk management, fixed income securities, financial engineering, and monetary theory and policy. Our goal is to provide a solid theoretical foundation and a wealth of practical examples for the benchmark dataset.
@@
Appendix D Statistical Characterization of the FinEval-KR Dataset
A complete sample from the constructed dataset is shown in Figure 10. The sample size and its distribution for each subdiscipline in FinEval-KR Dataset is shown in Figure 11, where the subdiscipline categorization methodology refers to Zhang et al. (2023). The top-50 knowledge points in the dataset are shown in Figure 12. In each subdiscipline, the distribution of cognitive labels is shown in Figure 13.
@@ Appendix D Statistical Characterization of the FinEval-KR Dataset
Appendix E Prompt Template Adopted by the FinEval-KR Evaluation Framework
Figure 15 shows the prompt template used in Stage 1 where the evaluated model answers the questions in a free-form format. Figure 16 shows the prompt template used in Stage 2, in which the evaluated model re-answer the question with knowledge point augmented. Note that we have made every effort to ensure the consistency of the prompt structure used in Stage 1 and Stage 2 to avoid changes in reasoning performance due to structural variations. During the prompt design, we conduct multiple rounds of trials and adjustments, ultimately ensuring that changes in the model’s responses are solely related to the injected knowledge points.
Figure 14 shows the prompt templates for the judge model to generate review results and augmented review results based on the reference answers in Stage 1 and 2.
@@ Appendix E Prompt Template Adopted by the FinEval-KR Evaluation Framework
Appendix F Details of Experiment
Bias Challenges in Judge Model
@@ Appendix F Details of Experiment
  • Style Bias: This bias refers to the tendency of the judge LLM to give higher scores to content with a more appealing text style (e.g., clear structure, moderate length), even if the answers contain reasoning errors Wu and Aji (2025). To reduce the impact of this type of bias, we did not restrict the output format of the models being evaluated in Stage 1 and Stage 2, encouraging them to reason freely. Subsequently, we used methods such as regular expressions to unify the original output of each model into a format consistent with the reference answer. This processing method effectively reduced evaluation biases caused by differences in text style.
  • Cognitive Bias: This bias refers to the self-bias that LLMs may exhibit during the evaluation, i.e., a tendency to give higher scores to content they generated themselves, thereby affecting the fairness of the evaluation Koo et al. (2024). To avoid this type of cognitive bias, we excluded GPT-o1 from the scope of evaluated models in our experiment, as it had already been used in the dataset construction and validation. Furthermore, during the preliminary experiment, we tested whether Qwen2.5-72B_Instruct, used as the judging model, exhibited significant cognitive bias in the root cause localization and knowledge gap identification tasks. The experimental results showed that this model did not exhibit a significant self-bias tendency in the aforementioned two tasks. We believe this may be because these two tasks are more specific and objective compared to result comparison tasks without sub-task decomposition, and are further aided by the model introspection prompting shown in Figure 14, which helps enhance the objectivity of Qwen2.5-72B_Instruct during the evaluation.
  • To prevent the judge model from the disturbance of “simple deception” Thakur et al. (2024), we filter out meaningless content in the generated answers, such as isolated affirmative words like “yes” or “of course”, ensuring that the evaluation focuses on the substantive reasoning process rather than superficial linguistic features.

### v3

@@ 3.2 Results Analysis
4 FinEval-KR
In this section, we present a Financial domain Evaluation framework for assessing Knowledge and Reasoning abilities (FinEval-KR). We first describe the methodology used to construct the evaluation dataset. Next, we introduce a multi-stage evaluation framework that performs root cause analysis of reasoning errors via knowledge-augmented question answering, enabling a decoupled assessment of a model’s knowledge mastery and reasoning ability. Finally, we define a series of evaluation metrics: a knowledge score based on domain knowledge coverage; a reasoning score and a cognitive score which is based on Bloom’s taxonomy. This approach improves the interpretability of LLM evaluations in financial scenarios and provides clear directions for targeted model improvement.
4.1 Benchmark Dataset Construction
Corpus Collection
To ensure the dataset is both authoritative and relevant, we selected nine canonical textbooks from major financial disciplines. These sources provide a comprehensive and up-to-date overview of modern finance. This process yielded a financial corpus totaling 8,460 pages. A detailed list of the textbooks and the rationale for their selection is available in Appendix C.
Question Generation
We generate financial problems from the obtained corpus using a two-stage process. First, we use a custom-designed prompt to instruct OpenAI o1, a state-of-the-art model at the time of our research that was specially enhanced for reasoning capabilities, to create a computational question based on a given text segment. The prompt’s design ensures the question meets predefined standards (see Figure 6 in the appendix for details). Second, we subject each candidate question to a three-step automated validation: it is checked for logical coherence, consistency with the source material, and overall quality. Questions that fail any check are discarded, ensuring the high fidelity of the final dataset. The details of validation please refer to Appendix D.
Answer Generation
We generate and validate the ground-truth answers using a structured three-stage pipeline: (1) Unconstrained generation: We first prompt OpenAI o1 to solve each problem, guided by the prompt shown in Figure 7. Crucially, we impose no initial format constraints on the output, a strategy designed to capture the model’s most natural and diverse problem-solving pathways. (2) Standardized formatting: The resulting raw solutions are then systematically parsed and reformatted according to a predefined template to ensure consistent and uniform presentation across the dataset. (3) Rigorous validation: Finally, each formatted answer undergoes a triple-validation protocol, which mirrors the question validation process (see Appendix D for details).
Dataset Annotation
For each question, we use OpenAI o1 guided by the prompt in Figure 8 to identify all requisite knowledge points. These include, but are not limited to, core financial concepts, regulatory frameworks and mathematical operations.
For each step in the answer, we annotate the corresponding cognitive level using Bloom’s Taxonomy. To mitigate the inherent subjectivity of this task, we developed a constrained, keyword-driven methodology. We guide the OpenAI o1 using a predefined set of keywords strongly associated with each cognitive level, derived from  anderson2001taxonomy (see Figure 9 for prompt template). This process maps each reasoning step to one or more levels: Remembering, Understanding, Applying, Analyzing, and Evaluating. We intentionally exclude the Creating level. This decision ensures objectivity, as our dataset comprises problems with determinate solutions, a principle that aligns with standardized financial certification exams.
Dataset Characteristics and Exemplary Samples
The final evaluation set contains a total of 9,782 question-answer pairs and their associated annotations. The questions and answers in the dataset are verified by human experts, and a sample check showed that the accuracy of the dataset is above 90%. Further details on the human annotators, the quality control mechanisms, and the validation process are provided in the Appendix J. An exemplary sample from the constructed dataset is given in Figure 1 on a yellow background. (see Figure 10 for the complete example). The complete statistical characterization of the dataset is detailed in Appendix E.
4.2 Evaluation Framework
As shown in Figure 1, the FinEval-KR framework comprises three evaluation stages. First, a model attempts a problem, and an LLM-as-a-judge identifies any errors, extracting the specific knowledge points needed for a correct solution. In the second stage, we provide the model with this missing knowledge and have it re-answer the question. The final stage performs a comparative analysis for attributing the initial error to knowledge deficiency or reasoning ability deficiency. This mechanism allows for the independent quantification of a model’s knowledge and reasoning capabilities.
4.2.1 Stage 1: Question Answering
Unconstrained Solution Generation
In this initial stage, the model under evaluation generates a solution without any format constraints. This design is crucial for preventing judgment errors based on superficial format-matching. It ensures, for example, that a logically sound reasoning path is not unfairly penalized simply for deviating from the step-order of the reference answer. The prompt for this stage is detailed in Appendix F, Figure 15.
Structured Judgment and Error Analysis
Next, we employ an LLM-as-a-judge to analyze the generated solution against a reference answer. To ensure the judgment is objective and reproducible, the judge is guided by a highly structured, Chain-of-Thought (CoT) prompt that enforces a rigorous step-by-step analysis (see Figure 14). Additionally, we address the potential bias challenge of the judge model in Appendix G.
If an error is detected, the judge’s output, termed the review result, pinpoints the first incorrect step, identifies its root cause, and lists the corresponding knowledge deficiencies (an example is shown in Figure 2). If the final answer is correct, we consider the entire reasoning process valid. This assumption is grounded in the novelty and multi-step complexity of our dataset, which makes correct answers via guessing or exploiting artifacts (the “Clever Hans” effect) highly improbable.
4.2.2 Stage 2: Knowledge-augmented Answering
If the evaluated model makes a error in Stage 1, the framework proceeds to this second stage. Here, we provide the model with the exact knowledge points that the judge identified as missing in Stage 1. This knowledge is integrated into a new prompt (see Appendix F, Figure 16), instructing the model to re-attempt the problem. The core purpose of this stage is to isolate the reasoning variable. By explicitly providing the necessary knowledge – a prerequisite for correct reasoning as validated in our preliminary study – we can now assess if the model can reason correctly when its knowledge gaps are filled. The judge then re-evaluates the new solution using the same protocol as in Stage 1. The outcome of this assessment is termed the augmented review result.
4.2.3 Stage 3: Error Diagnosis
The final stage performs a comparative analysis between the outcomes of Stage 1 and Stage 2 to determine the root cause of the initial error. Our approach is grounded in the principle that LLM’s preference to external information reveals its internal knowledge gaps (Wu2024How). Specifically, the judge model compares the review result with the augmented review result, following the principles below to determine the root cause of reasoning errors:
  • Knowledge Deficiency. If the augmented review result shows that the model reasons correctly in Stage 2, or the erroneous step occurs later than in Stage 1. This indicates that the evaluated model preferred the augmented knowledge in the second stage. This proves that the initial error is caused by knowledge deficiency.
  • Reasoning Ability Deficiency. If the evaluated model still makes a reasoning error in Stage 2, and the erroneous step is consistent with Stage 1, this indicates that the evaluated model still preferred its internal prior knowledge. This proves that the initial error is rooted in poor reasoning ability.
Through this attribution method, our framework successfully decouples a model’s knowledge and reasoning abilities from its overall task performance.
4.2.4 Evaluation Metrics
We propose three core metrics: knowledge score, reasoning score, and cognitive score, while also retaining accuracy to measure the model’s overall task performance.
Knowledge Score (KS)
$KS=1-\frac{\left|\bigcup_{i=1}^{M}K^{\prime}_{i}\right|}{\left|\bigcup_{i=1}^{N}K_{i}\right|},$ (1)
Reasoning Score (RS)
$RS=\frac{\sum_{i=1}^{N}\mathbb{I}(a_{i}=a_{i}^{\text{ref}})}{N-\sum_{i=1}^{N}\mathbb{I}(a_{i}\neq a_{i}^{\text{ref}}\land r(a_{i})=\mathrm{K})},$ (2)
where $a_{i}$ is the answer of the evaluated model for the $i$-th question, and $a_{i}^{\text{ref}}$ is the corresponding reference answer. Thus, the numerator of Eq. (2) is the total number of questions with correct reasoning. $r(a_{i})$ represents the root cause for the erroneous result $a_{i}$, which can be either $\mathrm{K}$ or $\mathrm{R}$, representing knowledge and reasoning ability deficiency, respectively. In addition, since each reasoning step in the dataset is annotated with a cognitive label, the reasoning ability deficit can be further subdivided into the ability deficit at a certain cognitive level, i.e., $\mathrm{R}^{j}$, $j\in\{1,2,3,4,5\}$, where $j$ is the index of the level in Bloom’s taxonomy, and $1,2,3,4,5$ denote remembering, understanding, applying, analyzing, and evaluating respectively. The denominator in Eq. (2) is the total number of questions whose reasoning errors are not attributed to knowledge deficiencies.
Cognitive Score (CS)
Furthermore, to explore the level of cognition exhibited by LLMs when solving reasoning tasks, we expanded a series of fine-grained metrics drawing on RS. The $j$-th cognitive level score of the LLM is defined as,
$CS_{j}=RS\times(1$- (3)
$\alpha_{j}\times\frac{\sum_{i=1}^{N}\mathbb{I}(r(a_{i})=\mathrm{R}^{j})}{N-\sum_{i=1}^{N}\mathbb{I}(a_{i}\neq a_{i}^{\text{ref}}\land r(a_{i})=\mathrm{K})}).$
$\alpha_{j}\in(0,1)$ is a penalty coefficient, and it is designed to have negative correlation with the cognitive level $j$. This weighting scheme is designed to heavily penalize errors made by the model during lower-level cognitive reasoning steps. In our experiments, we empirically set $\alpha_{1},\cdots,\alpha_{5}$ to a linearly decreasing sequence 0.9, 0.8, 0.7, 0.6 and 0.5.
Accuracy (Acc)
Finally, the success rate for all reasoning tasks is defined as,
$Acc=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(a_{i}=a_{i}^{\text{ref}}).$ (4)
5 Experiments
This section first validates the FinEval-KR framework’s effectiveness in identifying potential knowledge weaknesses and performing root cause analysis. Subsequently, we selected a range of LLMs, and evaluated them using the proposed FinEval-KR framework and our open-sourced dataset.
5.1 Alignment Analysis of Judge Model in FinEval-KR
The core of FinEval-KR lies in the design of the judge model for identifying initial reasoning step, recalling potential knowledge weaknesses and pinpointing root causes, which is fundamental for achieving decoupled evaluation. To this end, we design a set of comparative experiments to assess the alignment between the judge model and human evaluation.
The methods participating in the comparison include: (a) Direct Prompting: We use advanced models like OpenAI o1, directly prompting them to execute knowledge identification and root cause localization tasks. (b) Task Decomposition: Decompose the above two tasks into multiple subtasks and clearly define the logical dependencies between these subtasks (c) Ours: Based on task decomposition, this method requires the model to self-reflect and explicitly output reasoning processes step-by-step before generating the final conclusion (i.e., adding the <Inner Thought> as shown in Figure 14).
Given that our benchmark dataset is composed in Chinese, we select Qwen2.5-72B_Instruct, which is specifically optimized through extensive pre-training on Chinese corpora, as backbone LLM in the latter two methods. We use the following three metrics to evaluate the performance of these methods: (a) Accuracy of error identification: This refers to the proportion of correctly localized initial reasoning step in the Stage 1. (b) Accuracy of recalled knowledge points: The percentage of missing knowledge correctly recalled in Stage 1. (c) Accuracy of error attribution : The proportion of correctly attributed reasoning errors to knowledge or reasoning ability deficits.
Additionally, we discuss the limitations of adopting Qwen2.5-72B_Instruct as the judge model in the Limitations section.
Methods
Error
Identification
Knowledge
Recall
Error
Attribution
Direct
Prompting
0.56 0.24 0.20
Task
Decomposition
0.85 0.60 0.53
Ours 0.92 0.94 0.93
Table 2: Accuracy for three evaluation tasks.
Model/Metrics Acc KS RS
CS1
CS2
CS3
CS4
CS5
Qwen2.5-14B_Instruct 0.5473 0.8490 0.6863 0.6547 0.6603 0.3893 0.6863 0.6820
QwQ-32B-preview 0.7380 0.9073 0.8627 0.8450 0.8503 0.6987 0.8510 0.8597
DeepSeek-v3 0.8270 0.9427 0.9077 0.8963 0.8993 0.7963 0.8943 0.9057
DeepSeek-R1 0.8700 0.9517 0.9347 0.9377 0.9397 0.8810 0.9380 0.9433
Doubao-pro-32k 0.7825 0.9195 0.8750 0.8560 0.8600 0.7340 0.8565 0.8720
Moonshot-v1-128k 0.4533 0.8340 0.6020 0.5620 0.5670 0.2763 0.5653 0.5973
Ernie-bot-4.0 0.5733 0.8627 0.7053 0.6680 0.6753 0.4383 0.6847 0.6927
Qwen-max-latest 0.6467 0.8797 0.7733 0.7507 0.7547 0.5340 0.7440 0.7703
GPT-3.5-turbo 0.2830 0.7527 0.3973 0.3527 0.3603 0.0900 0.3893 0.3970
GPT-4o 0.6853 0.9020 0.8067 0.7847 0.7890 0.5930 0.7870 0.8030
GPT-4.1 0.8263 0.9520 0.9063 0.8957 0.8977 0.7890 0.8927 0.9050
o1-mini 0.7503 0.8997 0.8453 0.8340 0.8363 0.6983 0.8477 0.8450
o3-mini 0.8207 0.9260 0.9070 0.9047 0.9073 0.8127 0.9023 0.9120
Gemini-2.5-pro 0.8750 0.9627 0.9233 0.9123 0.9163 0.8403 0.9050 0.9120
Gemini-2.5-flash 0.8440 0.9540 0.9203 0.9103 0.9133 0.8307 0.9100 0.9177
Claude-3.7-sonnet 0.7923 0.9390 0.8823 0.8663 0.8703 0.7433 0.8653 0.8803
Xuanyuan-FinX1-preview 0.5890 0.8687 0.7323 0.7063 0.7130 0.4610 0.7323 0.7300
Fin-R1-7B 0.4153 0.7510 0.5570 0.5190 0.5277 0.2170 0.5570 0.5527
5.2 Evaluation
For our evaluation, we select 18 leading and representative LLMs, referencing prominent leaderboards like the Chatbot Arena [^2] https://openlm.ai/chatbot-arena/. This selection spans a diverse range of models, including open-source and closed-source systems, models of varying parameter scales, and different architectures such as dense and Mixture-of-Experts (MoE) (see Appendix H for details).
In our analysis, we focus on the relative performance rankings (i.e., performance tiers) of these models rather than their absolute scores. This approach is designed to ensure the robustness of our findings. While an LLM-as-a-judge may have inherent systematic biases, such biases have a smaller impact on the relative ordering of models than on their absolute scores.
5.3 Results and Core Findings
Table 3 lists the average metrics for all 18 models from three independent runs, conducted with a model temperature of 1. The complete results, including standard deviations, are presented in Table 7.
The analysis of the model performance echelons and discussion of the results are detailed in in Appendix I. In summary, the comprehensive analysis of all evaluation metrics identifies the current Tier 1 models as DeepSeek-R1, Gemini-2.5-pro and Gemini-2.5-flash. These models typically have parameters in excess of a hundred billion and use the MoE architecture to optimize computational resources. Furthermore, they specifically optimize reasoning capabilities through methods like reinforcement learning, and demonstrate outstanding performance in knowledge coverage and the completeness of reasoning paths.
Bottleneck in Knowledge Applying Abilities
Our analysis reveals that it is reasoning and specific cognitive skills, not merely knowledge, that truly drive performance in advanced LLMs. While the Knowledge Score (KS) and Reasoning Score (RS) are positively correlated, KS scores converge among top-tier models, indicating that sheer knowledge is no longer the primary performance bottleneck. Instead, RS shows a strong correlation with accuracy, establishing reasoning ability as crucial for success. A deeper cognitive analysis pinpoints the ability to apply knowledge (CS3) as the critical differentiator, evidenced by a sharp drop in this metric, which directly degrades their reasoning and accuracy.
Crucially, this weakness is not confined to lower-tier models. Even top-tier LLMs exhibit a significant deficit in applying knowledge (CS3) compared to their abilities in analyzing (CS4) or evaluating (CS5). For instance, GPT-4.1 scores 0.7890 in CS3 versus 0.8927 in CS4. This universal shortcoming underscores a fundamental limitation of current models: a profound difficulty in transferring theoretical knowledge to practical, real-world application.
The Dilemma of Financial LLMs
Our evaluation establishes Xuanyuan-FinX1-preview as the leading specialized financial LLM, consistently outperforming counterparts like Fin-R1-7B. However, a more critical finding emerges when comparing it to state-of-the-art general LLMs. Despite its domain leadership, Xuanyuan-FinX1-preview exhibits a significant performance gap of above 20%, a deficit that spans across financial complex reasoning, and higher-order cognitive skills. We attribute this gap to the superior generalization capabilities of leading general LLMs, which are developed through pre-training on vast, multi-domain datasets and benefit from more rapid iteration cycles. This advantage allows them to achieve strong performance even in financial fields, highlighting the limitations of current financial LLMs in terms of data diversity and development velocity.
Consequently, we predict a dual-track future for developing high-performance financial LLMs. The first track involves building upon state-of-the-art general foundation models to leverage their vast world knowledge and robust generalization. The second requires employing advanced fine-tuning techniques to distill and align these models for specialized financial reasoning, moving beyond a simple reliance on domain data.
6 Conclusion
Limitations
A potential limitation of this study lies in the choice of the judge model. Our primary experiments were conducted using Qwen-2.5-72B_Instruct, which represented the state-of-the-art among publicly available models with strong Chinese support during our experimental phase in late 2024. With the rapid evolution of large language models, even more capable reasoning models like DeepSeek-R1 have since emerged.
To investigate the impact of this evolution, we performed a small-scale evaluation using DeepSeek-R1 as the judge model. The results revealed a clear performance-efficiency trade-off: while DeepSeek-R1 yielded a marginal accuracy improvement of approximately 3%, it nearly doubled the inference time, posing significant challenges for large-scale evaluation. Crucially, we found that although the absolute scores of the evaluated models slightly increased, their relative rankings and the performance gaps between them remained highly consistent. Since our research focuses on the comparative performance of different methods, this consistency confirms the robustness of our conclusions.
Our future work will focus on enhancing the framework’s robustness by incorporating reasoning LLMs and more diverse evaluation paradigms. Specifically, we plan to employ multiple, heterogeneous models for dataset generation and implement a cross-validated, multi-judge evaluation pipeline to minimize potential biases, raise the evaluation ceiling, and bolster the benchmark’s overall reliability.
Acknowledgement
We thank the anonymous reviewers for their helpful suggestions. This work was supported by Ant Group Research Fund.
Appendix A The Dataset and Prompts Used in Preliminary Experiments
@@ Appendix C Data Sources for the FinEval-KR Dataset
To ensure our benchmark is both authoritative and comprehensive, we constructed the source corpus from nine classic textbooks in modern finance. This selection provides extensive coverage across key subfields, including corporate finance, investments, financial markets, risk management, and monetary policy. These foundational texts supply a rich combination of core theoretical principles and practical case studies, forming a robust basis for evaluating financial knowledge and reasoning.
We processed the corpus using a three-stage pipeline: extraction, cleaning, and standardization. (1) Extraction: We used OCR to convert all text and mathematical equations from the source materials into a machine-readable Markdown format. (2) Cleaning: We then manually curated the extracted content, removing non-essential sections (e.g., prefaces, appendices) and performing quality assurance checks. (3) Standardization: Finally, we transformed the cleaned content into a structured format suitable for automated processing. This rigorous process ensures the final dataset is of high quality, integrity, and utility.
@@
Appendix D Validation in Question and Answer Generation
For both the question and answer generation phases, we adopted a three-stage verification process. The verification focus for each stage is detailed in Table 4 and Table 5, respectively.
Stage Aspect Criteria
Logical Validation Clarity and Completeness Check (1) Is the question description clear and unambiguous? (2) Is there any misuse of terminology? (3) Are all necessary conditions and data for calculation provided?
Plausibility Check (1) Are numerical values (e.g., interest rates, returns, prices) within a plausible range? (2) Is the scenario self-contradictory or unrealistic?
Solvability Check Assuming the data is complete and plausible, does the question have a deterministic solution that can be calculated using financial models?
Consistency Assessment Relevance Check Do the core concepts in the question align with the title or keywords of the corresponding textbook chapter?
Final Validation Samples that fail any of the above checks are marked as "unqualified" and removed from the final dataset.
Table 4: Validation in question generation stage.
Stages Aspect Criteria
Logical Validation Formula/Model Selection Check (1) Is the selected formula a standard method for this type of problem? (2) Does the variable required by the question match the variable solved for in the answer?
Parameter Substitution Check Do the numerical values from the question correctly correspond to the variables in the formula during calculation?
Consistency Assessment Calculation Validation Execute the calculation code output by OpenAI o1 to verify the answer’s correctness.
Final Validation Samples that fail any of the above checks are marked as "unqualified" and removed from the final dataset.
Table 5: Validation in answer generation stage.
Appendix E Statistical Characterization of the FinEval-KR Dataset
A complete sample from the constructed dataset is shown in Figure 10. The sample size and its distribution for each subdiscipline in FinEval-KR Dataset is shown in Figure 11, where the subdiscipline categorization methodology refers to zhang2023fineval. The top-50 knowledge points in the dataset are shown in Figure 12. In each subdiscipline, the distribution of cognitive labels is shown in Figure 13.
@@ Appendix E Statistical Characterization of the FinEval-KR Dataset
Appendix F Prompt Template Adopted by the FinEval-KR Evaluation Framework
Figure 15 shows the prompt template used in Stage 1 where the evaluated model answers the questions in a free-form format. Figure 16 shows the prompt template used in Stage 2, in which the evaluated model re-answer the question with knowledge point augmented. Figure 14 shows the prompt templates for the judge model to generate review results and augmented review results based on the reference answers in Stage 1 and 2.
@@ Appendix F Prompt Template Adopted by the FinEval-KR Evaluation Framework
Appendix G Bias Challenges in Judge Model
@@
  • Style Bias: This bias refers to the tendency of the judge LLM to give higher scores to content with a more appealing text style (e.g., clear structure, moderate length), even if the answers contain reasoning errors wu2025style. To reduce the impact of this type of bias, we did not restrict the output format of the models being evaluated in Stage 1 and Stage 2, encouraging them to reason freely. Subsequently, we used methods such as regular expressions to unify the original output of each model into a format consistent with the reference answer. This processing method effectively reduced evaluation biases caused by differences in text style.
  • Cognitive Bias: This bias refers to the self-bias that LLMs may exhibit during the evaluation, i.e., a tendency to give higher scores to content they generated themselves, thereby affecting the fairness of the evaluation koo2024benchmarking. To avoid this type of cognitive bias, we excluded OpenAI o1 from the scope of evaluated models in our experiment, as it had already been used in the dataset construction and validation. Furthermore, during the preliminary experiment, we tested whether Qwen2.5-72B_Instruct, used as the judging model, exhibited significant cognitive bias in the root cause localization and knowledge gap identification tasks. The experimental results showed that this model did not exhibit a significant self-bias tendency in the aforementioned two tasks. We believe this may be because these two tasks are more specific and objective compared to result comparison tasks without sub-task decomposition, and are further aided by the model introspection prompting shown in Figure 14, which helps enhance the objectivity of Qwen2.5-72B_Instruct during the evaluation.
  • To prevent the judge model from the disturbance of “simple deception” thakur2024judging, we filter out meaningless content in the generated answers, such as isolated affirmative words like “yes” or “of course”, ensuring that the evaluation focuses on the substantive reasoning process rather than superficial linguistic features.
Appendix H Details of Experiment